a. c. 25.5. Using the tables and partial fractions, find the inverse Laplace transform for each of the following: 75+5 5-1 b. Please only do d,h,i

Thanks!
Answer #1

2. 121. (A) 38246842크 598 38168f27 A BRAC روسی و + 49 3s + 6A +2+ = A894 BB 3s + GA+27 - S (A+B) -4 28 +9A 사려 3 At = 3 3+B =

(क) 23}+382+28+27 - F(8) (3279) (871) 2.8333+23+27 ABAB (8²79 ) (3²711 8379 BUD 3+1 28+38++25+27 Asztastbe%B7C$*+9cstes:49 (3

Terve 2 Cos (31) + 3 sint A 13 1 St $13 3+7 cil 68² + 62 84 92 Frola (54116.1241084211 68² + 62 8+ 92 (SH/C+3)(at) 6.844 62.3